OpenStudy (anonymous):

Solve 5x^2 = -30x - 65. a. x = -3 ± 2 b. x = -3 ± 4 c. x = -6 ± 2 d. x = -6 ± 4

OpenStudy (anonymous):

Move everything on one side, 5x^2 + 30x + 65 = 0 Apply Quadratic Formula, \[\frac{ -B \pm \sqrt{B^2-4AC} }{ 2A }\]

OpenStudy (anonymous):

As stee stated, you use the quadratic formula, but you're going to have complex numbers. Once you plug in your numbers into the Quadratic formula, and split it you'll end with the following. \[\frac{ -6-\sqrt{-16} }{ 2 } and \frac{ -6+\sqrt{-16} }{ 2 }\] You can't take the square of a negative number, but if you remember \[i=\sqrt{-1}\] \[\frac{ -6-4 i }{ 2 } and \frac{ -6+4i }{ 2 }==Simplify==>-3-2i and -3+2i\]

OpenStudy (shamil98):

Actually, you can take the gcf of 5 out and then use the quadratic formula. 5x^2 + 30x + 65 = 0 5(x^2 + 6x + 13) = 0 x^2 + 6x + 13 = 0 and then use the quadratic formula.

Directrix (directrix):

@lowcard2 Please check to determine if you correctly posted the answer options. There seems to be something missing.

OpenStudy (anonymous):

a. x = -3 ± 2 i